Find the domain of :


f(x)=(16-x)C{Combination}(2x-1)+(20-3X)P(PERMUTATIONS)(4X-5)

Grade:12

1 Answers

Chetan Mandayam Nayakar
13 years ago

2x-1 is a whole number. therefore 2x-1=n, x=(n+1)/2, similarly 4x-5=m,x=(m+5)/4,(n+1)/2=(m+5)/4,thus m=2n-3.

16-x≥2x-1,x≤17/3,n≤31/3,  20-3x≥4x-5, x≤25/7,m≤65/7

therefore m=1,3,5,7,9  and the domain of the function is 3/2,2,5/2/3,7/2

'Win exciting gifts by answering the questions on Discussion Forum.'